# CMS Results at the t threshold

### v1

@@ 1 Introduction
Top-Antitop quark pair production and its subsequent decays offer a promising opportunity to research beyond the standard model (BSM) theories. In particular, models with particles featuring Yukawa-like couplings exhibit a big discovery potential here as the coupling is largest due to the high top quark mass. Many models that fit these criteria include extended Higgs sectors, such as 2 Higgs Doublet Models (2HDM), which also occur in the context of Supersymmetry. A key feature of these models is - among other new particles - the presence of heavy pseudoscalar (A) and scalar Higgs boson (H). For large parts of the extended Higgs parameter space these two particles dominantly decay to a top quark-antiquark pair. In addition to BSM signatures, for masses close to 2*, there is a potential contribution from bound states, predicted in earlier calculations [Kiyo:2008bv, Sumino:2010bv, Ju:2020otc]. The top quark bound state effects are expected to form a broad peak below the t threshold and exhibit dominantly pseudoscalar behavior at the LHC. They have not previously been observed. With these two signatures being very similar, a search for both BSM and SM scenarios is based on the same methodologies. This work presents such a search [CMS-HIG-22-013], using the full Run 2 dataset of the CMS detector [Chatrchyan:2008zzk, CMS:2023gfb] with an integrated luminosity of 138 fb-1. This work superseedes the previous CMS search for A and H bosons with only 35.9 fb-1 of data [CMS:2019pzc]. The results consider both the semileptonic () decay channel and the dileptonic () decay channel. For the threshold bound state results in [CMS:2025kzt], exclusively the dileptonic decays are considered.

### v2

@@ 1 Introduction
Top-Antitop quark pair production and its subsequent decays offer a promising opportunity to research beyond the standard model (BSM) theories. In particular, models with particles featuring Yukawa-like couplings exhibit a big discovery potential here as the coupling is largest due to the high top quark mass. Many models that fit these criteria include extended Higgs sectors, such as 2 Higgs Doublet Models (2HDM), which also occur in the context of Supersymmetry. A key feature of these models is - among other new particles - the presence of heavy pseudoscalar (A) and scalar Higgs boson (H). For large parts of the extended Higgs parameter space these two particles dominantly decay to a top quark-antiquark pair. In addition to BSM signatures, for masses close to 2*, there is a potential contribution from bound states, predicted in earlier calculations [Kiyo:2008bv, Sumino:2010bv, Ju:2020otc, Fadin:1987wz, Fadin:1990wx]. The top quark bound state effects are expected to form a broad peak below the t threshold and exhibit dominantly pseudoscalar behavior at the LHC. They have not previously been observed. With these two signatures being very similar, a search for both BSM and SM scenarios is based on the same methodologies. This work presents such a search [CMS-HIG-22-013], using the full Run 2 dataset of the CMS detector [Chatrchyan:2008zzk, CMS:2023gfb] with an integrated luminosity of 138 fb-1. This work superseedes the previous CMS search for A and H bosons with only 35.9 fb-1 of data [CMS:2019pzc]. The results consider both the semileptonic () decay channel and the dileptonic () decay channel. For the threshold bound state results in [CMS:2025kzt], exclusively the dileptonic decays are considered.